1401 BCD Encodings in ASCII and Tables for Addressing above 1K

The following pages have tables of:

• Pierce Primary (A chain) BCD encoding,

• Pierce Alternative (H chain) BCD encoding, and

• Traditional SimH BCD encoding.

Each table includes a Secret Decoder Ring for addressing above 1k.

Page 2 has six copies of Pierce Primary encoding; page 3 has six copies of Pierce Alternative encoding; page 4 has six copies of SimH encoding.

If you print pages 2 and 3 duplex and page 4 simplex, you can laminate the sheets and cut each one into six pieces, giving shirt-pocket size reference cards for the three encodings.

Page 5 has two copies of all three encodings, but even if cut in half each set of tables is a bit too large for a shirt-pocket reference.

Page 6 has three copies of the Pierce encodings, while page 7 has three copies of the SimH encoding, with space for one table below each. If you print pages 6 and 7 duplex and laminate the sheet, you can cut it into three shirt-pocket size reference cards with Pierce encodings one one side and SimH encodings on the other.

Notes: Zero (0) is represented internally as C82 but punched as 0. Record mark ( ⊧ on 1403, | in Pierce encodings, ’ in SimH encoding) is represented internally as A82 but is punched as 0-2-8 — because 0 can’t simultaneously be used for A and 82. The A bit alone (ˆ ) has no input punch on 1401 without a free RPQ (which number I’ve forgotten). With the RPQ or on 1410 it’s 2-8. It might punch as 0.

— Van Snyder 23 March 2010

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Pierce Primary (A chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , % ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < }

Pierce Alternative (H chain) Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 = ’ : > {
0  A  4 5 6  7 ∧ / S T U V W X Y Z | , ( ∼ \ ”
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 + A B C D E F G H I ? . ) [ < }

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Notes: Zero (0) is represented internally as C82 but punched as 0. Record mark ( ⊧ on 1403, | in Pierce encodings, ’ in SimH encoding) is represented internally as A82 but is punched as 0-2-8 — because 0 can’t simultaneously be used for A and 82. The A bit alone (ˆ ) has no input punch on 1401 without a free RPQ (which number I’ve forgotten). With the RPQ or on 1410 it’s 2-8. It might punch as 0.

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Notes: Zero (0) is represented internally as C82 but punched as 0. Record mark ( ⊧ on 1403, | in Pierce encodings, ’ in SimH encoding) is represented internally as A82 but is punched as 0-2-8 — because 0 can’t simultaneously be used for A and 82. The A bit alone (ˆ ) has no input punch on 1401 without a free RPQ (which number I’ve forgotten). With the RPQ or on 1410 it’s 2-8. It might punch as 0.

Traditional SimH BCD Encoding

Card Core  Core  1 2 3 4 5 6 7 8 9 0 Card Row Punch
Zone Units  Hundreds Zone  Core Digit Bits  1 2 3 4 5 6 7
Row Zone none A B AB 0 1 2 3 4 5 6 7 8 8 8 8 8 8 8 8
none none  0 1 2  3  1 2 3 4 5 6 7 8 9 0 # @ : > (
0  A  4 5 6  7 ∧ / S T U V W X Y Z ’ , % = \ +
11  B  8 9 10 11 - J K L M N O P Q R ! $ * ] ;
12  AB  12 13 14 15 & A B C D E F G H I ? . ) [ < ”

Notes: Zero (0) is represented internally as C82 but punched as 0. Record mark ( ⊧ on 1403, | in Pierce encodings, ’ in SimH encoding) is represented internally as A82 but is punched as 0-2-8 — because 0 can’t simultaneously be used for A and 82. The A bit alone (ˆ ) has no input punch on 1401 without a free RPQ (which number I’ve forgotten). With the RPQ or on 1410 it’s 2-8. It might punch as 0.